answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What has a head a tail and is round?

Judging by what you are saying, if it has a curly tail it obviously a pig.

When 2 dice are thrown what is its probability of accuring 7?

1/6

To have a sum of 7 we need

1 and 6

or

2 and 5

or

3 and 4

If we look at 1 and 6, we could get a 6 or a 1 first, the chance of

either one of those is 2/6 or 1/3

Now once we have that, we need an exact number for the second number.

If we rolled a 6 we need a 1. We have a 1/6 chance of that.

So we have (1/3)(1/6)=1/18

Now we can do this for each of the three ways we have to get 7, that is to say for (1,5), (2,5) and (3,4).

So we add 1/18+1/18+1/18=3/18 or 1/6

How do you get the probability of dice rolls?

A normal die is a six faced cube, so the probability of each face coming up is one in six and there are six possible outcomes. If you throw two dice, then there are 36 (6x6) possible outcomes, and you would simply divide the desired result set by the possible result set. For instance, the probability of throwing a 12 is 1 in 36, while the probability of throwing an 11 is 2 in 36 or 1 in 18 because there are two ways to throw the 11. Similarly, the probability of throwing a 7 is 6 in 36 or 1 in 6, because there are 6 different ways to throw a 7. Continuing with more and more dice, it is a simple matter of counting and arithmetic.

What is the value of 1789-1797 George Washington two head coin?

It's a novelty coin with no collectible value. Just like the 2 headed Kennedy half dollars

What is the intersection of the assessed probability?

Leaders can supervise compliance with hazard controls during a mission

What major events occur in telophase?

Telophase occur in both mitosis and meiosis. During telophase nuclear membrane and nuclei reforms and in meiosis it further goes for meiosis 2 which is same as mitosis

What does complement mean in probability?

The complement of an event occurring is that it does not occur.

How many spades appear on a ace deck of cards?

Three spades appear on the ace in a standard deck of cards.

Does a king and jack beat a king and 4 high cards?

Yes the king and jack would win over a king and 4 high cards

In Statistics what is a quantity that measures the variation of a population or sample relative to its mean?

Inter-quartile range, other percentile ranges, mean absolute variation, variance, standard error, standard deviation are all possible measures.

What are the chance of you having a girl if your family only has boys?

well, generally the chances are 50/50. You should probably know however, if you don't know already, that the chromosome determining the child's sex comes only from the father, so check his family line. If there are a lot of men in it, well, there is a good chance there will be boys. If there are a lot of girls, there is a good chance of having girls. You really can't tell though until you actually have the child! If you simply MUST have a girl, you can do alternative methods such as artificial insemination, where doctors isolate your partners male sperm from the female sperm, and put the selected genders of sperms with several of your eggs that they have removed from your body. once they have conceived, they will place the cells back into your body where they will grow and form a baby of your selected sex. this costs a lot of money though.

Differentiate between Worst case analysis and Monte Carlo analysis?

A complete Six Sigma Mentoring System. This is going to be the industry standard in driving Lean Six Sigma projects. Voice guided with instructions how and where to. Has Continuous training as it walks you through the DMAIC methodology, does statistical analysis, reports and tracks the whole project. No limits to amount of projects, users or capabilities. We will custom tailor to your companies operating applications. This is 1/5 the price of the leading industry software with everything they do not offer in training and analysis. Call 713-436-6941 A demonstration is a must.

Probability of picking a winning 3 digit lottery number?

One out of a thousand


^^^This is not quite true. If you must match them in order, then this is correct. However, if order does not matter (i.e. 4-5-6 is the same as 6-4-5), then the probability increases.

The reason is that the probability is no longer 1/10*1/10*1/10, since this would assume that all three are unrelated. If order doesn't matter, then 4-5-6 = 6-4-5 = 5-4-6 = ... and so on, and all of these win. Thus, the answer is (all the different ways you can write the 3 digit number)/1000.

What is the relationship between tarot and probability theory?

Those who are skeptics would say Tarot has everything to do with random occurences. Those who believe know that the messages received are from the spirit realm and are not connected. I do tell my clients that some things I see can only be predicted by probabilities because we have free will. For instance, if a question is asked about the future of a relationship, I can tell my client that it is probable that this relationship will lead to marriage. I cannot be 100% sure because either party could have a change of heart at any time. This is where free will comes in. In other types of questions, I can be more precise. In the end, it comes down to our perception and belief system. I have clients that I have been reading for many years and they report that I am highly accurate. If my answers to them were based on random occurences, my accuracy rate would be much lower and my clients would not return to me year after year.

If you have 15 cards with different shapes If the first card you draw is a square and the probability is 1 in 5 What is the probability of next card something other than a square?

1 in 5 is a square so there are 3 squares and 12 non-squares. If the square is not replaced, and the next card is drawn at random the probability of a non-square is 12/14.

A man takes a step forward with probability 0.4 and backwards with probability 0.6 Find the probability that at the end of eleven steps he is just one step away from the starting point?

1. He takes 6 steps forward and 5 steps backward

2. He takes 5 steps forward and 6 steps backward

consider 1:

p1 = (11C6)(0.4)^6*(0.6)^5

for 2:

p2 = 11C5(0.4)^5*(0.6)^6

total probability = p1 + p2

Trending Questions
You are throwing a 8 sided die with the numbers 1-8 inclusive on the sides What is the probability for each outcome? What is the probability of drawing either a Jack or a 5 from a deck of cards? What are the chance of you having a girl if your family only has boys? What is the probability of drawing a black card with 2 from a deck of 52 cards? When two dice are rolled find the probability of getting a sum less than 4 or greater than 9? What is the or rule in probability? If you roll a fair six number cube what is the probability of rolling two numbers whose sum is two? What is the probability that parents that have 2 boys and one girl will have a girl next? You have 6 vertices and 5 faces? What is the probability of rolling a 5 on the first roll or a 4 on the second roll? How many ways can the letters abcdef be arranged? When you roll three number cubes what is the probability that all three numbers will match? A single card is drawn from a deck of cards what is the probability of the card being a seven or a heart? What is the probability that you will pick an odd number or a number greater than 13 from 1 to 20? In a bag of 25 marble's 7 are blue 4 are green 8 are red and the rest are yellow what is the ratio of blue marble's to yellow marble's? How many possible outcomes would there be if a coin is tossed? How many marbles does Paolo have in the bag Paolo has less than 40 marbles in a bag and splits them among 7 friends and has 2 left over? How many ways can you get a sum of 11 when 2 dice are thrown? What does total probability equal? What are the chances of a card not being a heart in a deck of cards?